OpenStudy (anonymous):

Evaluate the expression when p = –4 and q = 2. 20 + 6q ÷ p A. 8 B. –17 C. –8 D. 17

OpenStudy (michele_laino):

if we replace \(p=-4,q=2\), we get: \(20+6\cdot(2):(-4 )\)

OpenStudy (michele_laino):

now, we have to do the multiplication and division first, from left to right

OpenStudy (michele_laino):

so, what is \(6 \cdot 2=...\)

OpenStudy (anonymous):

12

OpenStudy (michele_laino):

then, what is \(12:(-4)=...\)

OpenStudy (anonymous):

3

OpenStudy (michele_laino):

are you sure?

OpenStudy (michele_laino):

we have one minus sign

OpenStudy (anonymous):

-3

OpenStudy (michele_laino):

finally, what is: \(20-3=...\)

OpenStudy (anonymous):

17

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

thank you

OpenStudy (michele_laino):

:)